how much +ve and – ve charge is  there in cup of water .suppose 250 has a water in cup

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
 
Here quantity of water in cup is 250 ml
That means its weight is 250 gm
Number of moles = 250/water molar weight
                            = 250/18 = 13.88
Number of molecules = 13.88 x6.022x10^23
                                   = 8.36x 10^24
Total proton = total electron = 10
Number of proton in water = 8.36 x10^25
Charge = 1.6 * 10^-19 C
Total ppositive charge = 13.37 x 10^6 C = 13.3MC
Hence total positive and negative = 13.37 x 10^6 C = 13.3MC hence the overall charge is zero.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)